Prove that every non-trivial subgroup of a cyclic group has finite index. Hence prove that (Q,+) is not cyclic.


1
Expert's answer
2021-05-07T11:39:42-0400

Given, G is a cylic group.


Therefore, G=<a>.


And H is subgroup of G.


H=<a^i>.


Index,


"\dfrac{G}{H} ={a\hat{j}+<a\hat{i}>}"


If j>i then by division algorithm, j=ir+s


Then "a\hat{j}+<a\hat{i}>=a\hat{s}+<a\hat{i}>"


So, "\dfrac{G}{H}={a\hat{s}+<a\hat{i}>, 0<=s<i}"


"O(\dfrac{G}{H})" = finite =index of H.


For( Q,+), choose H="<\dfrac{1}{2}>"


Clearly "O(\dfrac{G}{H})=" inifinte. 


So, G=(Q,+) is not cyclic.